On complete combustion, 0.246 g of an organic compound gave 0.198 g of CO2 and 0.1014 g of H2O. What is the ratio of carbon and hydrogen atoms in the compound?

1
Expert's answer
2022-03-16T14:28:03-0400

All of the carbon from the hydrocarbon is converted to carbon dioxide:


"n(C)=n(CO_2 )=\dfrac{mass}{molar\ mass}=\dfrac{0.198\ g}{12+2*16\ g/mol}=0.0045\ mol"


All of the hydrogen from the hydrocarbon is converted to water (each water molecule contains two hydrogen atoms):


"n(H)=2*n(H_2O )=2*\dfrac{mass}{molar\ mass}=2*\dfrac{0.1014\ g}{2*1+16\ g/mol}=0.0113\ mol"


Carbon and hydrogen ratio:


"H:C=0.0113:0.0045=2.5:1=5:2"